OpenStudy (anonymous):

Sally travels by car from one city to another. She drives for 22.0 min at 52.0 km/h, 46.0 min at 42.0 km/h, and 47.0 min at 27.0 km/h, and she spends 5.0 min eating lunch and buying gas. (a) Determine the average speed for the trip.km\h (b) Determine the total distance traveled. km

OpenStudy (anonymous):

shane thanks for the help on the last one i figured that one out

OpenStudy (shane_b):

If you want I can check your answers once you've calculated these

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

for average i got 1.008 for total distants i got 121.0

OpenStudy (anonymous):

\[52.0+42.0+27.0=121.0\]

OpenStudy (anonymous):

\[22.0+46.0+47.0=115+5=120\]

OpenStudy (shane_b):

I must be tired...I goofed up a symbol in my initial post. Here it is again: \[speed=\frac{distance}{time}\]a) To calculate the average speed you take the total distance traveled and divide that by the total time it took to cover that distance. b) The total distance traveled will be the sum of the distances for each leg of her trip. Having said that, please redo the calcs.

OpenStudy (shane_b):

Since speed = distance/time, distance = speed x time. That's how you calculate the distance for each leg.

OpenStudy (anonymous):

i added the distances together and i added the time together so all i have to do is \[121.0\120 \right?\]

OpenStudy (shane_b):

No. I'll show you how to calculate the distance of the first leg: 22min @ 52 km/hr Therefore the distance for that leg will be:\[(\frac{22}{60}h)*52km/h=19.067km\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

u forgot to say i have to turn the min into hours

OpenStudy (shane_b):

You always have to convert units if they are different :)

OpenStudy (anonymous):

thats the step i was mising

OpenStudy (anonymous):

now i calculated all 3 legs now what

OpenStudy (shane_b):

What distances did you get for each leg? You just need to sum them up to get a total and then divide that by the total time of 120 mins (2 hrs).

OpenStudy (anonymous):

19.07,32.2,21.15 there right webassign confirmed it now for the average speed

OpenStudy (anonymous):

72.42 total distance

OpenStudy (shane_b):

Right....so the average speed would be:\[Speed_{avg}=\frac{72.42km}{2hr}=?\]

OpenStudy (anonymous):

why 2 hours?

OpenStudy (shane_b):

120 minutes = 2 hours. The original speeds were given in km/h so I would put the answer in km/h.

OpenStudy (anonymous):

u included that extra 5 minutes at the end?

OpenStudy (shane_b):

Yes...remember you added them up (and I checked): 22.0+46.0+47.0+5.0=120

OpenStudy (anonymous):

then is it 36.21?

OpenStudy (shane_b):

Yep

OpenStudy (anonymous):

nice! got a 76 on the webassign so far ur a great teacher

OpenStudy (shane_b):

I probably could have explained it better but you got it :)

OpenStudy (anonymous):

can i ask the next question in here

OpenStudy (shane_b):

No guarantees but maybe I can help with it

OpenStudy (anonymous):

okay

OpenStudy (shane_b):

You should post it as a separate question just in case...so everyone else can see it

OpenStudy (anonymous):

ok